Direction: Solve the following problems stated below. Show your solutions.



1. Find the resistance of a 190m titanium wire with cross sectional area of 0.55mm^2. The resistivity of titanium is 420Ωm.



2. A piece of number 10 aluminum wire has a diameter of 2.59mm. How many meters of number 10 aluminum wire is needed to give a resistance of 1Ω? The resistivity of the aluminum is 2.8x10^-8Ωm.

1
Expert's answer
2022-02-07T15:28:15-0500


1. Resistance:


"R=\dfrac{\rho L}{A}=14.51\text k\Omega"


2. The length:


"L=\dfrac{RA}{\rho}=\dfrac{R\pi d^2}{4\rho}=188\text{ m}."